# On the completeness problem of homogeneous Lorentz 3-dimensional spaces, Part II

### v1

@@ Abstract.
We consider 3-dimensional solvable Lie groups given by a semi-direct product $\mathbb{R}\ltimes_{A}\mathbb{R}^{2}$, where $A$ is a nonzero real $2\times 2$ diagonalizable matrix. The completeness problem has been addressed in the literature for two limiting cases: the unimodular Lie group $\mathrm{Sol}$ and the homothetic Lie group $\mathrm{Ho}$. In this work, we solve the completeness problem of left-invariant Lorentzian metrics on the remaining Lie groups of the given form. Moreover, we show that every 3-dimensional non-unimodular Lie group has an incomplete left-invariant metric.
@@ 1. Introduction
      and $\mathfrak{z}$ and $\mathfrak{d}$ are orthogonal, q is complete and all integral curves of the geodesic field are unbounded.

### v2

# The completeness problem on 3-dimensional non-unimodular Lie groups

@@ Abstract.
We consider the completeness problem for left-invariant Lorentzian metrics on 3-dimensional non-unimodular Lie groups, all of which have Lie algebra of the form $\mathbb{R}\ltimes_{A}\mathbb{R}^{2}$, where $A$ is a real $2\times 2$ matrix with nonzero trace. The case where $A$ is not diagonalizable over $\mathbb{C}$ was addressed in previous work by the authors, and the limiting case where $A$ is a scalar multiple of the identity is also known from the literature. In this paper, we determine all geodesically (in)complete left-invariant Lorentzian metrics for all other cases where $A$ is diagonalizable over $\mathbb{R}$. Additionally, we show that, when $A$ is diagonalizable over $\mathbb{C}$ but not over $\mathbb{R}$, there exists at least one incomplete metric. As a consequence of prior work and our results, we obtain that every 3-dimensional non-unimodular Lie group admits an incomplete left-invariant Lorentzian metric.
@@ 1. Introduction
      and $\mathfrak{z}$ and $\mathfrak{d}$ are orthogonal, q is complete and all non-stationary integral curves of the geodesic field are unbounded.